OpenStudy (anonymous):

solve the equation: (z^3)-1=0

OpenStudy (anonymous):

\[z ^{3}-1=0\]

OpenStudy (anonymous):

1

OpenStudy (anonymous):

including imaginary roots

OpenStudy (anonymous):

cube roots of 1 are what you want

OpenStudy (agreene):

\[\large z=1~~~z=\sqrt[3]{-1} ~~~ z=(-1)^{2/3}\]

OpenStudy (anonymous):

1 is one of them others are taken by dividing up the unit circle in to 3 equal parts you will get \[-\frac{1}{2}+\frac{\sqrt{3}}{2}i\] and its conjugate

OpenStudy (anonymous):

before we got reset, i tried to write solve \[(z-1)(z^2+z+1)=0\] quadratic formula will give the second two solutions